# Three Fundamental Questions in Modern Infinite-Domain Constraint Satisfaction

### v2

@@ A.1. Omitted definitions
Suppose that $\mathrm{age}(\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}})$ has the RP. Then clearly also $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP. Let $\mathfrak{X},\mathfrak{Y}\in\mathrm{age}(\mathfrak{B})$ be arbitrary. We transform them into $\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ by letting $E$ interpret as the diagonal relation and $<$ as the empty relation. Since $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP, there exists a Ramsey witness $\mathfrak{W}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$. We set $\mathfrak{W}\coloneqq(\mathfrak{W}^{\prime}/E^{\mathfrak{W}^{\prime}})|^{\sigma}$. Let $f\colon\binom{\mathfrak{W}}{\mathfrak{X}}\rightarrow[2]$ be arbitrary. We define $f^{\prime}\colon\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}\rightarrow[2]$ as follows. For an embedding $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$, we set $f^{\prime}(e^{\prime})\coloneqq f(q_{E}\circ e^{\prime})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$. We have that $q_{E}\circ e^{\prime}\in\binom{\mathfrak{W}}{\mathfrak{X}}$ because $E$ interprets as the diagonal relation in $\mathfrak{X}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$. Since $\mathfrak{W}^{\prime}$ is a Ramsey witness for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$, there exists $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{Y}^{\prime}}$ such that $f^{\prime}$ is constant on $\bigl{\{}e^{\prime}\circ u^{\prime}\;|\;u^{\prime}\in\binom{\mathfrak{Y}^{\prime}}{\mathfrak{X}^{\prime}}\bigr{\}}\subseteq\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$. Since $E$ interprets as the diagonal relation in $\mathfrak{Y}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$, the map $e\coloneqq q_{E}\circ e^{\prime}$ is contained in $\binom{\mathfrak{W}}{\mathfrak{Y}}$. By construction, $f$ is constant on $\bigl{\{}e\circ u\;|\;u\in\binom{\mathfrak{Y}}{\mathfrak{X}}\bigr{\}}\subseteq\binom{\mathfrak{W}}{\mathfrak{X}}$.
@@
Proof of item 4. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=((\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}})|^{\tau}$. Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
For the other direction, we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4]. Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. Since $\mathfrak{A}$ and $\mathfrak{C}$ are homomorphically equivalent, $\mathfrak{C}$ also does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 1.3 in [7], there exist $\alpha,\beta,s\in\mathrm{Pol}(\mathfrak{C})$ witnessing the pseudo-Siggers identity. Consider the structure $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$. By item 1, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is an $\omega$-categorical model-complete core. Note that we can lift the homomorphic equivalence between $\mathfrak{A}$ and $\mathfrak{C}$ to a homomorphic equivalence between $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ by setting, for any homomorphism $h$ between $\mathfrak{A}$ and $\mathfrak{C}$ and irrespectively of its direction, $h^{\ast}(x,y)\coloneqq(e(x,y),h(y))$, where $e$ is an arbitrary injection from the Cartesian product of $\mathbb{Q}$ with the domain of $h$ into $\mathbb{Q}$. The inclusion of the injection in the first coordinate ensures that $h^{\ast}$ preserves $I_{4}$ and $\neq$ (because they are preserved by all injective operations). Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ are homomorphically equivalent, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. Next, we lift the pseudo-Siggers identity from $\mathrm{Pol}(\mathfrak{C})$ to $\mathrm{Pol}((\mathbb{Q}\wr\mathfrak{C})_{\neq})$. Let $e$ be an arbitrary injection from ×$\mathbb{Q}\times C$ into $\mathbb{Q}$. The structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. Hence, by Theorem 1.3 in [7] (since pp-constructions are more general than pp-interpretations with parameters), there are $\alpha^{\prime},\beta^{\prime},s^{\prime}\in\mathrm{Pol}(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity. We set $\alpha^{\ast}(x,y)\coloneqq(\alpha^{\prime}(x),\alpha(y))$, $\beta^{\ast}(x,y)\coloneqq(\beta^{\prime}(x),\beta(y))$, and
@@ Appendix B $\omega$-categorical cheeses for PCSPs
for all $s_{1},\dots,s_{n}\in S$. Suppose, on the contrary, that $f$ is not cyclic. Then there are $s_{1},\dots,s_{n}\in S$ such that $f(s_{1},\dots,s_{n})\neq f(s_{2},\dots,s_{n},s_{1})$. Let $<$ denote the linear order of $\mathfrak{B}$ and set $\tilde{s_{i}}\coloneqq(s_{i},\dots,s_{n},s_{1},\dots,s_{i-1})$. By assumption, we have $f(\tilde{s_{1}})\neq f(\tilde{s_{2}})$. Without loss of generality, $f(\tilde{s_{1}})<f(\tilde{s_{2}})$. Since $\alpha_{1},\alpha_{2}$ preserve $<$ and its complement, we have
$f(\tilde{s_{1}})<f(\tilde{s_{2}})\Leftrightarrow$ $\enspace\alpha_{1}(f(\tilde{s_{1}}))<\alpha_{1}(f(\tilde{s_{2}}))$
$\Leftrightarrow$ $\enspace\alpha_{2}(f(\tilde{s_{2}}))<\alpha_{2}(f(\tilde{s_{3}}))$
$\Leftrightarrow$ $\enspace f(\tilde{s_{2}})<f(\tilde{s_{3}})$
@@
$\Leftrightarrow$ $\enspace f(\tilde{s_{n}})<f(\tilde{s_{1}}),$
@@
The next lemma essentially states that $\mathcal{I}(\mathfrak{X})\coloneqq{\mathfrak{X}}^{[d]}$ is a reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. We will, however, use it in the more general setting of PCSPs.
@@
Let $\mathfrak{A}$ be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. Suppose that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. Then, for every finite structure $\mathfrak{X}$ and every homomorphism $h\colon{\mathfrak{X}}^{[d]}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$, there exists a homomorphism $g\colon\mathfrak{X}\to\mathfrak{A}$ such that $h(\bar{x})=g(\bar{x})_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ for every $\bar{x}\in X^{d}$.
@@
[…]
@@
$x\sim y:\Leftrightarrow$ $\exists\bar{z}=(z_{1},\dots,z_{d})\in X^{d}\,\exists i_{1},i_{2}\in[d]$
[…]
@@
$x\sim y\Leftrightarrow$ $\forall\bar{z}=(z_{1},\dots,z_{d})\in X^{d}\,\forall i_{1},i_{2}\in[d]\colon$
@@
$\bar{y_{1}}\in R^{\mathfrak{X}^{\prime}}$ (1)&(2)${\Longleftrightarrow}h(\bar{z_{1}})\in R_{\mathrm{id}}^{\,\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$
@@
[…]
@@
Let $\mathfrak{A}$ be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$ and let $\mathfrak{S}$ be a finite substructure of $\mathfrak{A}$. Suppose that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. If $\Sigma$ is a height-1 condition that is satisfied in $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$, then $\Sigma$ is satisfiable in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$.
@@
It is not hard to show that ${(\mathfrak{S}^{n})}^{[d]}\simeq({\mathfrak{S}}^{[d]})^{n}$ (cf. [57, Corollary 22]) for all $d,n\in\mathbb{N}$. Hence, every homomorphism $h\colon({\mathfrak{S}}^{[d]})^{n}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ is also a homomorphism from ${(\mathfrak{S}^{n})}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Suppose that $h$ is a function symbol appearing in $\Sigma$ that is assigned an element of $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. By Lemma B.2, there is a homomorphism $g\colon\mathfrak{S}^{n}\to\mathfrak{A}$ such that
@@
for all $\bar{s_{1}},\dots,\bar{s_{d}}\in S^{n}$. For each function symbol in $\Sigma$, we fix such a homomorphism. Suppose now that $h_{1},h_{2}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ witness the satisfaction of the identity
@@
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have ${\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Let $S\subseteq A$ with $|S|\geq 3$. Suppose that there is a finite cheese $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\coloneqq h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $n$-ary $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}).$ By Lemma B.7 there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. By Lemma B.1, this operation is cyclic. Take pairwise distinct elements $s_{1},s_{2},s_{3}\in S$. We now show that the cyclicity of $g$ contradicts the fact that it preserves $\neq$. To this end, we consider the following three cases:
@@
In all three cases, $g$ does not preserve $\neq$. Hence, the original assumption that such $\mathfrak{D}$ exists was wrong, and consequently $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is not finitely tractable. ∎

### v3

@@ A.1. Omitted definitions
Suppose that $\mathrm{age}(\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}})$ has the RP. Then clearly also $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP. Let $\mathfrak{X},\mathfrak{Y}\in\mathrm{age}(\mathfrak{B})$ be arbitrary. We transform them into $\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ by letting $E$ interpret as the diagonal relation and << as the empty relation. Since $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP, there exists a Ramsey witness $\mathfrak{W}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$. We set $\mathfrak{W}\coloneqq(\mathfrak{W}^{\prime}/E^{\mathfrak{W}^{\prime}})|^{\sigma}$. Let $f\colon\binom{\mathfrak{W}}{\mathfrak{X}}\rightarrow[2]$ be arbitrary. We define $f^{\prime}\colon\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}\rightarrow[2]$ as follows. For an embedding $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$, we set $f^{\prime}(e^{\prime})\coloneqq f(q_{E}\circ e^{\prime})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$. We have that $q_{E}\circ e^{\prime}\in\binom{\mathfrak{W}}{\mathfrak{X}}$ because $E$ interprets as the diagonal relation in $\mathfrak{X}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$. Since $\mathfrak{W}^{\prime}$ is a Ramsey witness for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$, there exists $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{Y}^{\prime}}$ such that $f^{\prime}$ is constant on $\bigl{\{}e^{\prime}\circ u^{\prime}\;|\;u^{\prime}\in\binom{\mathfrak{Y}^{\prime}}{\mathfrak{X}^{\prime}}\bigr{\}}\subseteq\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$. Since $E$ interprets as the diagonal relation in $\mathfrak{Y}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$, the map $e\coloneqq q_{E}\circ e^{\prime}$ is contained in $\binom{\mathfrak{W}}{\mathfrak{Y}}$. By construction, $f$ is constant on $\bigl{\{}e\circ u\;|\;u\in\binom{\mathfrak{Y}}{\mathfrak{X}}\bigr{\}}\subseteq\binom{\mathfrak{W}}{\mathfrak{X}}$.
@@
Proof of item 4. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=((\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}})|^{\tau}$. Thus, if $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
For the other direction, we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4]. Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. Since $\mathfrak{A}$ and $\mathfrak{C}$ are homomorphically equivalent, $\mathfrak{C}$ also does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 1.3 in [7], there exist $\alpha,\beta,s\in\mathrm{Pol}(\mathfrak{C})$ witnessing the pseudo-Siggers identity. Consider the structure $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$. By item 1, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is an $\omega$-categorical model-complete core. Note that we can lift the homomorphic equivalence between $\mathfrak{A}$ and $\mathfrak{C}$ to a homomorphic equivalence between $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ by setting, for any homomorphism $h$ between $\mathfrak{A}$ and $\mathfrak{C}$ and irrespectively of its direction, $h^{\ast}(x,y)\coloneqq(e(x,y),h(y))$, where $e$ is an arbitrary injection from the Cartesian product of $\mathbb{Q}$ with the domain of $h$ into $\mathbb{Q}$. The inclusion of the injection in the first coordinate ensures that $h^{\ast}$ preserves $I_{4}$ and $\neq$ (because they are preserved by all injective operations). Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ are homomorphically equivalent, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. Next, we lift the pseudo-Siggers identity from $\mathrm{Pol}(\mathfrak{C})$ to $\mathrm{Pol}((\mathbb{Q}\wr\mathfrak{C})_{\neq})$. Let $e$ be an arbitrary injection from $\mathbb{Q}\times C$ into $\mathbb{Q}$. The structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. Hence, by Theorem 1.3 in [7] (since pp-constructions are more general than pp-interpretations with parameters), there are $\alpha^{\prime},\beta^{\prime},s^{\prime}\in\mathrm{Pol}(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity. We set $\alpha^{\ast}(x,y)\coloneqq(\alpha^{\prime}(x),\alpha(y))$, $\beta^{\ast}(x,y)\coloneqq(\beta^{\prime}(x),\beta(y))$, and
@@ Appendix B $\omega$-categorical cheeses for PCSPs
[…]
$\Leftrightarrow$ $)\displaystyle\enspace\alpha_{2}(f(\tilde{s_{2}}))<\alpha_{2}(f(\tilde{s_{3}}$))
[…]
$\vdots\$
[…]
@@
[…]
@@
(cf. Section 5.5 in [68]) Let $\mathfrak{A}$ be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. Suppose that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. Then, for every finite structure $\mathfrak{X}$ and every homomorphism $h\colon{\mathfrak{X}}^{[d]}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$, there exists a homomorphism $g\colon\mathfrak{X}\to\mathfrak{A}$ such that $h(\bar{x})=g(\bar{x})_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ for every $\bar{x}\in X^{d}$.
@@
$R^{\mathfrak{X}^{\prime}}\coloneqq\{(x_{\iota(1)},\dots,x_{\iota(k)})\in X^{k}\mid$ $\text{ there exists }\iota\colon[k]\to[d]\text{ and }(x_{1},\dots,x_{d})\in X^{d}$
$\text{ such that }h(x_{1},\dots,x_{d})\in R_{\iota}^{\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}\}$. (1)
@@
$x\sim y:\Leftrightarrow\$$\exists\bar{z}=(z_{1},\dots,z_{d})\in X^{d}\,\exists i_{1},i_{2}\in[d]$
$\textnormal{ with }x=z_{i_{1}},y=z_{i_{2}},\textnormal{ and }\forall\bar{a}=(a_{1},\dots,a_{d})\in h(\bar{z})\colon a_{i_{1}}=a_{i_{2}}.$
@@
$x\sim y\Leftrightarrow\$$\forall\bar{z}=(z_{1},\dots,z_{d})\in X^{d}\,\forall i_{1},i_{2}\in[d]\colon$
@@
$\ \overset{\text{def.}}{\Longleftrightarrow}\ \exists\bar{a}\in h(\bar{z_{1}})\cap R_{\mathrm{id}}^{{\mathfrak{B}}^{[d]}}$
$\ \overset{\text{def.}}{\Longleftrightarrow}\ \exists\bar{a}\in h(\bar{z_{1}})\cap R_{(i\mapsto k+1)}^{{\mathfrak{B}}^{[d]}}$
$\ \overset{\text{def.}}{\Longleftrightarrow}\ h(\bar{z_{1}})\in R_{(i\mapsto k+1)}^{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$
$\quad\ \vdots$
$\ \overset{\text{def.}}{\Longleftrightarrow}\ h(\bar{z_{2}})\in R_{\mathrm{id}}^{\,\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]